Ships A and B leave port together. For the next two hours, ship A travels at 25.0 mph in a direction 25.0 ∘ west of north while the ship B travels 20.0 ∘ east of north at 40.0 mph .

PART A:

What is the distance between the two ships two hours after they depart?

Express your answer with the appropriate units.

Part B:

What is the speed of ship A as seen by ship B?

Express your answer with the appropriate units.
Answer #1

Solution)

Ship A

x = 25 x sin(335) = -10.56 mph

y = 25 x cos(335) = 22.65 mph

Ship B

x = 40 x sin(20) = 13.68 mph

y = 40 x cos(20) = 37.58 mph

Now,

B - A:

x = 13.68 - (-10.56) = 24.24 mph

y = 37.58 - 22.65 = 14.93 mph

Relative speed = sqrt ( x^2 + y^2) = sqrt (24.24^2 + 14.93^2) = 28.46 mph

Distance covered after 2 h = 28.46*2 = 56.93 miles

So, speed of Ship A as seen by B = 28.46 mph (Ans)

========

Good luck!:)